Project NPV

H. Cochran, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2.15 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2.23 million in annual sales, with costs of $1.25 million. Assume the tax rate is 23 percent and the required return on the project is 14 percent. What is the project’s NPV? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.
Answer #1

Annual depreciation=(Cost-Residual value)/Useful Life = (2,150,000/3 ) = $716666.667

OCF = (Sales-Costs)(1-tax rate) + Tax savings on Annual depreciation = (2,230,000-1,250,000)(1-0.23) + (0.23*716666.667) = $919433(Approx).


answered by: gavin